. (3 pts) A 5.00 gram sample of an oxide of lead PbxOy contains 4.33 g of lead. Determine the simplest formula for the compound.
Answer #1

First calculate % compossition

% of Pb

5.00 gm of sample contain 4.33 gm lead

% if lead = gm of lead X 100 / gm of sample

% of lead = 4.33 X 100 / 5.00 = 86.6 %

% of lead = 86.6 %

% of O

% of O = 100 - % of lead = 100 - 86.6 = 13.4 %

% of O = 13.4 %

now calculate atomic ratio by dividing % compossition by atomic mass

atomic ratio for lead = 86.6 / 207.2 = 0.42

atomic ratio for O = 13.4 / 16 = 0.84

now calculate simplest ratio by dividing atomic ratio by smallest atomic ratio

simplest ratio for lead = 0.42/0.42 = 1

simplest ratio for O = 0.84/0.42 = 2

Simplest formula for compound is PbO2